OpenStudy (anonymous):

Can someone please help me simplify this logarithm: 4log(3) 1/3 + 2log(27) 9 1/3 and 9 are the arguments.

OpenStudy (lgbasallote):

hint: \[1) \; 4\log_3 (\frac 13) \implies 4\log_3 (3^{-1}) \implies 4(-1) \implies -4\] \[2) \; 2 \log_{27} (9) \implies 2\log_{27} (27^{2/3}) \implies 2 (\frac{2}{3}) \implies \frac 43\] does that help?

OpenStudy (anonymous):

Yes thank you you are amazing . But how do I make both of the bases equal??

OpenStudy (lgbasallote):

which base?

OpenStudy (lgbasallote):

do you mean what i did with \[2\log_{27} (27^{2/3})\]

OpenStudy (dumbcow):

you'd have to use change of base formula to do that..... log(b) A = log A/ log b

OpenStudy (lgbasallote):

im confused with this question....is this evaluating the logarithm? or is it condensing?

OpenStudy (anonymous):

Evaluating. It has an equal sign at the end and a "?"

OpenStudy (lgbasallote):

oh. then what do you mean by how to make the bases same?

OpenStudy (anonymous):

Well to solve this further and apply properties etc, don't I have to establish a common base?

OpenStudy (anonymous):

Lol, this is my summer assignment for Pre Calc and I completely forgot all the Algebra II.

OpenStudy (lgbasallote):

well...unless your condensing...you dont need to make the bases the same. you can just do what i did

OpenStudy (anonymous):

maybe I'm condensing I'm not sure. I just want to make this expression as simplified as possible and after what you did (thankyou) I got a step further but I don't know wht to do next.

OpenStudy (lgbasallote):

but...if you insist... \[\large 4\log_3 (\frac 13) \implies 4 \frac{\log_{27} (\frac 13)}{\log_{27} (3)} \implies 4 \frac{\log_{27} (3^{-1})}{\log_{27} (27^{1/3})} \implies 4 \frac{\log_{27} (27^{-1/3})}{\log_{27} (27^{1/3})}\] \[\large \implies 4(\frac{-1/3}{1/3} \implies 4(-1) \implies 4\] just a lengthy way to solve lol

OpenStudy (lgbasallote):

anyway.... what you do now is \[-4 + \frac 43\] then you have the answer

OpenStudy (anonymous):

Your phenomenal. I don't get it, but your fantastic. Thnk you for your time!! I will try to go over it. But you know, if you want to continue helping me, can you explain your steps? Hehe :P

OpenStudy (lgbasallote):

lol nevermind that second solution i did...just focus on the first one

OpenStudy (lgbasallote):

my first comment is the simplest way

OpenStudy (lgbasallote):

anyway i have to go now sorry

OpenStudy (anonymous):

Alright. I get it. Thanks <3 !!!

OpenStudy (lgbasallote):

welcome